OpenStudy (anonymous):

Y=X^3-1 Select integers for x,-3≤ x ≤ 3

OpenStudy (anonymous):

is the task to calculate the y value when x is between -3 and 3 ???

OpenStudy (anonymous):

it's to graph each equation between those two

OpenStudy (anonymous):

yeah... so you'll need the y-values to plot the x, y pairs... ???

OpenStudy (anonymous):

so calculate: (-3)^3 - 1 = ??? (-2)^3 - 1 = ??? (-1)^3 - 1 = ??? ( 0)^3 - 1 = ??? ( 1)^3 - 1 = ??? ( 2)^3 - 1 = ??? ( 3)^3 - 1 = ???

OpenStudy (anonymous):

oh yeah i was confused. i get it now thank you !